# A Photon Cloud Induced from an Axion Cloud

### v1

@@ IV The induced electric and magnetic fields
with respect to an Eulerian observer $n^{a}=\left(1/\alpha,-\beta^{i}/\alpha\right)$ whose worldline is orthogonal to the spacelike hypersurface.
Therefore we have
@@
In FIG. 2-4, we plot the components of induced electric and magnetic fields in cosine and sine sections respectively, in contrast with the background Wald case. In some panels with large scale like $100M$, we use the dotted circles to denote the radius $r_{\rm max}/M=2/\alpha_{1}^{2}$ of the maximum value of axion cloud, while in some panels with small scale like $5M$, we compare with the boundary of photon region (dotdashed curves) and event horizon (dashed circles). It is a good case to compare the distribution of the EMs (in field perspective) with the photon region (spherical photon orbits in particle perspective).
@@
Figure 2: The components of the induced electric fields are plotted as cosine part and sine part respectively. In all the figures we have set $a_{0}=0.8$ and $\alpha_{1}=0.2$.
@@
Figure 3: The components of the induced magnetic fields are plotted as cosine part and sine part respectively. In all the figures we have set $a_{0}=0.8$ and $\alpha_{1}=0.2$.
From the figures we can see that, the radial components $E^{r}_{\rm cos}$ and $E^{r}_{\rm sin}$ of the induced electric field both exhibit quadrupole moments with a parity violation ($P=-1$). In contrast, the radial component of electric field of the background Wald case is in dipole moments with a parity symmetry ($P=+1$). While the $\theta$ components $E^{\theta}_{\rm cos}$ and $E^{\theta}_{\rm sin}$ of the induced electric field both exhibit a parity symmetry ($P=+1$). In contrast, the $\theta$ component of electric field of the background Wald case is in quadrupole moments with a parity violation ($P=-1$). For the $\varphi$ components $E^{\varphi}_{\rm cos}$ and $E^{\varphi}_{\rm sin}$ of the induced electric field, they both show a parity violation symmetry ($P=-1$). Conversely, the symmetries for the components of magnetic fields are all opposite with those for electric fields. Besides, we can observe some evident coincides between the distribution of the induced electric and magnetic fields and the boundary of photon region.
Note that both electric and magnetic fields of the background Wald case do not contain $\varphi$ components, and they only distribute outside the ergoregion, while the induced EM field can live within the ergoregion. This can be understood as “no static observer is allowed inside the ergoregion”.
As we know, superradiance only happens for massive fields, where the mass term makes them gathering around the black hole and sustainably dissipate near the event horizon. The dissipation is a crucial ingredient for superradiance, in fact even for a rotating compact object without a horizon, superradiance can occur via other dissipation mechanism, like friction [10].
While for the axion-photon coupling case, due to the antisymmetry of $\epsilon^{\mu\nu\rho\sigma}$ there is never an effective mass term for the photons, though sometimes people call it “tachyonic-like instability” [20]. Also it has been mentioned that the time-dependent axion is a crucial ingredient to trigger this instability when $k_{a}\phi_{0}$ is larger than a critical value.
In our first-order study with an EM background, even for small couplings, the exponential growing happens for the EM field coupled to the axion field, as we have discussed in Sec. II. From our perspective, the reason is that the stationary axion cloud ties the EM field onto the nearby of the black hole and serves as a source continuously, which makes the EM field dissipate sustainably from the axion cloud to the photon region and finally to the event horizon. As it is shown in the figures, the induced EM field does not concentrate around the maximum value of axion cloud but its main distribution is inside the dotted circles. In other words, the most dense places for the photon cloud do not relate to the largest oscillating position (same radius $r_{\rm max}$) of the axion cloud but inside it.
@@
Figure 4: The components of the electric and magnetic fields are plotted for the background Wald case, where we have set $a_{0}=0.8$ and $\alpha_{1}=0.2$.
V Analysis of symmetries and parity violation
@@
Here in curved spacetime and classically, we obtain the same symmetry for the parity violation ($P=-1$). From (49), we have
where $m=1$ and
@@
all exhibit a sign change for $\theta\to\pi-\theta$, in agreement with the results in FIG. 5. Combined with no sign change for $\varphi\to\varphi+\pi$, we know that there is a parity violation $P=-1$ for $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$.
In addition, it shows that except for the evolving parts with a double frequency $2\omega$, there are also residential parts that are time-independent.
Figure 5: The components $M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos},~{}M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin},~{}M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$ of the parity violation term $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha_{1}=0.1,0.2$.
Similarly, the contraction of the EM field strength tensor $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ evolves as
where $m=1$ and
@@
Figure 6: The components $M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos},~{}M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin},~{}M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}$ of the contraction of the EM field strength tensor $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha_{1}=0.1,0.2$.
Moreover, we study the effects of the two parameters $a_{0},~{}\alpha_{1}$ and find that the spin parameter $a_{0}$ only has slight impacts on both the amplitude and the structure of the distribution for each components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$, while these quantities are sensitive to the fine structure constant $\alpha_{1}$. Therefore, in FIG. 5-6 we plot the figures for the components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ with fixed $a_{0}=0.8$ and different values of $\alpha_{1}$. From the figures we can see that, for all the components, the increase of $\alpha_{1}$ leads to a significant decrease in amplitude and a slight concentration on the distribution, which indicates that for fixed black hole mass $M$ the smaller axion mass $\mu_{1}$ (larger reduced Compton wavelength) makes the interaction with photons more efficient and stronger. Also, the slight concentration on the distribution should result from the concentration of the axion cloud with larger $\alpha_{1}$, where larger axion mass $\mu_{1}$ (smaller reduced Compton wavelength) makes the axion cloud concentrate. In general, the induced EM scalars $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ concentrate in the sine parts $M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}$ and $M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$.
@@
$E^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=-E^{r}_{\rm cos}\left(r,\theta\right)~{},\quad B^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=B^{r}_{\rm cos/sin}\left(r,\theta\right)~{},$ (66)
@@
which can be clearly seen in FIG. 2-3. We further have
@@
These transformations are equivalent to $\overset{\rightharpoonup}{E}\rightarrow-\overset{\rightharpoonup}{E}$ and $\overset{\rightharpoonup}{B}\rightarrow\overset{\rightharpoonup}{B}$ under $\overset{\rightharpoonup}{r}\rightarrow-\overset{\rightharpoonup}{r}$. Therefore the induced electric field still keep a parity symmetry with $P=-1$ as a vector, and the induced magnetic field keeps the parity symmetry with $P=+1$ as a pseudo-vector, same as the EM background.
In addition, we are interested in the strengths of the induced electric and magnetic fields that
which also works for $B^{i}_{(1)}B_{i}^{(1)}$.
We plot the strengths of electric and magnetic fields in different sections in FIG. 7, from which we can see that the strengths are mainly concentrated in $E^{2}_{\rm cos}$, $E^{2}_{\rm sin}$ and $B^{2}_{\rm sin}$, and for each panel there is a distinctive pattern of the distribution. It was mentioned in [31] that, when sources are present we can use the eigenfunctions of the angle-part equation to separate the master equation for $\varphi_{0}$ or $\varphi_{2}$ (their master equations without sources are separable) by expanding the source $T=\Sigma G(r)_{s}S_{lm}(\theta)e^{im\varphi}e^{-i\omega t}$. Remember that there are only two real degrees of freedom, i.e. once we get $\varphi_{0}$ or $\varphi_{2}$ then the other two complex quantities are decided. Therefore these featured patterns should come from the decomposition of the source, in our case it is the axion cloud but the expression of $T$ is so complicate to expand one by one.
Figure 7: The strengths of electric field and magnetic field $M^{2}E^{i}_{\rm cos}E^{\rm cos}_{i}$, $M^{2}E^{i}_{\rm sin}E^{\rm sin}_{i}$, $M^{2}E^{i}_{\rm cos}E^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm cos}_{i}$, $M^{2}B^{i}_{\rm sin}B^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm sin}_{i}$ are plotted with $a_{0}=0.8$ and $\alpha_{1}=0.2$.
VI Conclusions
In this work we studied the EM photon cloud sourced by an axion cloud through the axion-photon coupling around a Kerr black hole in the presence of a background EM field. In the first-order study of the coupling constant, we obtained an oscillating EM photon cloud with the same frequency as the axion cloud, which indicates that it exists even for very small couplings with no threshold. More intriguingly, this induced EM photon cloud is growing exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied.
In addition, we studied the induced electric and magnetic fields respectively, and found that their components exhibit very different symmetries in contrast with the EM background, which may indicate the existence of both an axion cloud and the axion-photon coupling. Furthermore, we analyzed the symmetries of most quantities in curved spacetime and confirmed the parity violation $P=-1$ of the $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ term. Besides, for near horizon region, we compared the distribution of the induced EM field with the boundary of photon region and found some overlap as expected.
Acknowledgements.
We thank Masahide Yamaguchi, Jin Sun, Yuhang Zhu and Cheng-Yong Zhang for beneficial discussions. This work is supported by IBS under the project code IBS-R018-D3.

### v4

@@ IV The induced electric and magnetic fields
with respect to an Eulerian observer $n^{a}=\left(1/\alpha,-\beta^{i}/\alpha\right)$ whose worldline is orthogonal to the spacelike hypersurface. Therefore we have
@@
The symmetries of these components are analyzed in Appendix C, consistent with the numerical results shown in FIG. 4-5.
@@
In FIG. 4-6, we plot the components of induced electric and magnetic fields in cosine and sine sections respectively, in contrast with the background Wald case. In some panels with large scale like $100M$, we use the dotted circles to denote the radius $r_{\rm max}/M=2/\alpha^{2}$ of the maximum value of axion cloud, while in some panels with small scale like $5M$, we compare with the boundary of photon region (dotdashed curves) and event horizon (dashed circles). It is a good case to compare the distribution of the EMs (in field perspective) with the photon region (spherical photon orbits in particle perspective).
@@
Figure 4: The dimensionless quantities $ME^{r}_{(1)}$, $M^{2}E^{\theta}_{(1)}$ and $M^{2}E^{\varphi}_{(1)}$ of the induced electric field components are plotted as cosine part and sine part respectively. In all the figures we have set $a_{0}=0.8$ and $\alpha=0.2$.
@@
Figure 5: The dimensionless quantities $MB^{r}_{(1)}$, $M^{2}B^{\theta}_{(1)}$ and $M^{2}B^{\varphi}_{(1)}$ of the induced magnetic field components are plotted as cosine part and sine part respectively. In all the figures we have set $a_{0}=0.8$ and $\alpha=0.2$.
@@
Figure 6: The dimensionless components $E^{r}_{\rm Wald}/B_{0}$, $ME^{\theta}_{\rm Wald}/B_{0}$, $B^{r}_{\rm Wald}/B_{0}$, $MB^{\theta}_{\rm Wald}/B_{0}$ of the electric and magnetic fields are plotted for the background Wald case, where we have set $a_{0}=0.8$.
From the figures we can see that, the radial components $E^{r}_{\rm cos}$ and $E^{r}_{\rm sin}$ of the induced electric field both exhibit quadrupole structures with a sign reversal across the equatorial plane. In contrast, the radial component of the electric field in the background Wald solution is in dipole structure that is symmetric about the equatorial plane. While the $\theta$-components $E^{\theta}_{\rm cos}$ and $E^{\theta}_{\rm sin}$ of the induced electric field both exhibit even symmetry with respect to the equatorial plane, the $\theta$-component of the electric field in the background Wald solution shows a quadrupole pattern with odd symmetry. In addition, the $\varphi$-components $E^{\varphi}_{\rm cos}$ and $E^{\varphi}_{\rm sin}$ of the induced electric field possess odd symmetry about the equatorial plane. Conversely, the magnetic field components have opposite symmetries to those of the electric field components. Besides, we can observe some evident coincides between the distribution of the induced electric and magnetic fields and the boundary of photon region.
Note that both electric and magnetic fields of the background Wald case do not contain $\varphi$-components, and they only distribute outside the ergoregion, while the induced EM field can live within the ergoregion. This can be understood as “no static observer is allowed inside the ergoregion”.
Moreover, we are interested in the strengths of the induced electric and magnetic fields that
which also works for $B^{i}_{(1)}B_{i}^{(1)}$.
Figure 7: The strengths of electric field and magnetic field $M^{2}E^{i}_{\rm cos}E^{\rm cos}_{i}$, $M^{2}E^{i}_{\rm sin}E^{\rm sin}_{i}$, $M^{2}E^{i}_{\rm cos}E^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm cos}_{i}$, $M^{2}B^{i}_{\rm sin}B^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm sin}_{i}$ are plotted with $a_{0}=0.8$ and $\alpha=0.2$.
We plot the strengths of the electric and magnetic fields in different sections in FIG. 7. It can be seen that the field strengths are mainly concentrated in $E^{2}_{\rm cos}$, $E^{2}_{\rm sin}$ and $B^{2}_{\rm sin}$, and that each panel exhibits a distinctive distribution pattern. As mentioned in [45], when sources are present, one can use the eigenfunctions of the angular equation to separate the master equation for $\varphi_{0}$ or $\varphi_{2}$ (their source-free master equations are separable) by expanding the source as $T=\Sigma G(r)_{s}S_{lm}(\theta)e^{im\varphi}e^{-i\omega t}$. Recall that there are only two real degrees of freedom; once $\varphi_{0}$ or $\varphi_{2}$ is determined, the other two complex quantities are fixed. Therefore, these featured patterns should come from the decomposition of the source. In our case, the source is the axion cloud, but the expression of $T$ is too complicate to expand term by term.
As it is well known, superradiance occurs only for massive fields, where the mass term appears in the effective potential as an asymptotic constant at spatial infinity, providing a naturally reflective boundary that traps waves with lower oscillation frequencies. In addition, the mass term localizes the field around the black hole and allows it to sustainably dissipate near the event horizon. The dissipation is a crucial ingredient for superradiance. In fact, even for a rotating compact object without a horizon, superradiance can occur through other dissipation mechanisms, such as friction [15].
In the case of axion-photon coupling, the antisymmetry of $\epsilon^{\mu\nu\rho\sigma}$ prevents the photons from acquiring an effective mass term for the photons; this is sometimes referred to as a “tachyonic-like instability” [28]. It has also been noted that a time-dependent axion is crucial for triggering this instability when $k_{a}\phi_{0}$ exceeds a critical value. Furthermore, as mentioned earlier, in the presence of an external magnetic field, a flat spacetime analysis can give the mixing matrices for both the axion-photon coupling and the dark photon-photon coupling cases [35]
$\mathcal{M}_{a\gamma}=\begin{pmatrix}-\omega_{p}^{2}/2\omega&B_{y}k_{a}\\ B_{y}k_{a}&-\mu_{a}^{2}/2\omega\end{pmatrix},\quad\mathcal{M}_{\gamma\gamma^{\prime}}=\frac{1}{2\omega}\begin{pmatrix}-\omega_{p}^{2}&\sin{\chi_{0}}\mu^{2}_{\gamma^{\prime}}\\ \sin{\chi_{0}}\mu^{2}_{\gamma^{\prime}}&-\mu_{\gamma^{\prime}}^{2}\end{pmatrix}$, (56)
where the presence of a magnetic field enables the axion-photon coupling to effectively act as a “mass term” in the mixing, whereas the dark photon-photon mixing matrix does not require an external magnetic field. Moreover, the flat spacetime analysis shows that in a homogeneous EM background, the dispersion relation is modified by the axion-photon coupling [33]
$\omega^{2}=\mathfrak{p}^{2}+\frac{1}{2}\mu^{2}\pm\frac{1}{2}\sqrt{\mu^{4}+16k_{a}^{2}E_{z}^{2}\mathfrak{p}^{2}}$~ (57)
where a sufficiently large electric field leads to $\omega^{2}<0$, corresponding to a positive imaginary part of $\omega$ and thus a growing mode (instability). This is very similar to the superradiance case. Particularly the axions and photons share the same frequency, although the instability is driven by a stronger electric field, unlike our case.
V Conclusions
In this work we studied the EM photon cloud sourced by an axion cloud through the axion-photon coupling around a Kerr black hole in the presence of a background EM field. At first order in the coupling constant, we obtained an oscillating EM photon cloud with the same frequency as the axion cloud, indicating that it can exist even for very small couplings, with no threshold for either the coupling strength or the EM background field amplitude.
More intriguingly, this induced EM photon cloud grows exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied within the perturbative framework. This instability is essentially inherited from the axion cloud, distinct from the instability that originates in the homogeneous part of the equation via the Mathieu equation, and also from that arising in the inhomogeneous (sourced) part of the EM field equation, where a sufficiently strong electric field is required.
The growth and spatial distribution of the induced EM field are thus primarily governed by the axion cloud and the EM background field, though with various featured patterns. Both the EM background and the axion cloud serve as sources driving the growth of the induced EM photon cloud, which is mainly concentrated near the black hole while dissipating sustainably from the axion cloud to the photon region and finally toward the event horizon. As shown in the figures, the induced EM field does not peak at the location of the maximum axion density; rather, its main distribution lies within the dotted circles. In other words, the densest regions of the photon cloud do not coincide with the radius of maximum axion oscillation (same as $r_{\rm max}$) but are located inside it.
In addition, we investigated the induced electric and magnetic fields respectively, and found that their components exhibit markedly different symmetries compared to those of the background EM field. As a result, the induced photon cloud forms an unstable bound configuration that emits EM waves to spatial infinity while being replenished by the axion cloud, providing a potential observational signature of both the presence of an axion cloud and axion-photon coupling. While a detailed comparison with black hole polarization images is beyond the scope of this work, the evolving EM structures induced by axion–photon coupling could, in principle, indirectly imprint subtle signatures in polarization observables—such as deviations from standard GRMHD polarization patterns or circular polarization—through their influence on plasma dynamics and radiative processes.
Furthermore, we analyzed the symmetries of most relevant quantities, which are consistent with the numerically results shown in the figures. Besides, in the near-horizon region, we compared the distribution of the induced EM field with the boundary of photon region and found partial overlap, as expected.
Acknowledgements.
We thank Masahide Yamaguchi, Jin Sun, Yuhang Zhu, Minxi He and Cheng-Yong Zhang for beneficial discussions. We thank the editor and the referees for their constructive comments and suggestions. We also thank Enrico Cannizzaro for his comments, which improved our interpretations. Z.-Y. Tang is supported by IBS under the project code IBS-R018-D3.
Appendix A Asymptotic behaviors of $f_{01M}$, $f_{11M}$ and $f_{\rm 21M}$
$f_{\rm 01M}\left(r_{\rm h0},\theta\right)=\frac{2\sqrt{2}a_{0}r_{h0}^{3}e^{-\alpha^{2}r_{\rm h0}/2}\sin{\theta}\sin{2\theta}}{\left(2\alpha r_{h0}-a_{0}-2i\sqrt{1-a_{0}^{2}}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~,$ (58)
$f_{\rm 11M}\left(r_{\rm h0},\theta\right)=\frac{2i\sqrt{1-a_{0}^{2}}r_{h0}^{3}e^{-\alpha^{2}r_{\rm h0}/2}\sin{2\theta}}{\left(a_{0}-2\alpha r_{h0}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~,$ (59)
$f_{\rm 21M}\left(r_{\rm h0},\theta\right)=0~,$ (60)
$f_{\rm 01M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha^{2}r_{0}/2}}{\sqrt{2}\alpha}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha^{2}r_{0}/2}\right)~,$ (61)
$f_{\rm 11M}\left(r_{0}\to\infty,\theta\right)=-\frac{(\alpha+3i)\sin{2\theta}e^{-\alpha^{2}r_{0}/2}}{4\alpha}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha^{2}r_{0}/2}\right)~,$ (62)
$f_{\rm 21M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha^{2}r_{0}/2}}{2\sqrt{2}\alpha}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha^{2}r_{0}/2}\right)~,$ (63)
In addition, it is clear that the asymptotic behaviors for $f_{\rm 01M}$ and $f_{\rm 11M}$ at event horizon $r_{\rm h0}$ are only valid when $\alpha\neq\frac{a_{0}}{2r_{\rm h0}}$, which is exactly the critical condition for whether superradiance occurs. In this critical case with $\alpha=\frac{a_{0}}{2r_{\rm h0}}$, the imaginary part $\omega_{I}$ disappears in the original field equation for the axion field, so that the whole frequency is not complex and the original boundary condition at event horizon is not valid, in which case another regular condition at origin should be applied. Here we do not go further on this case.
Appendix B Components of $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$
@@
The components of the parity violation term $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$
@@
all exhibit a sign change for $\theta\to\pi-\theta$, in agreement with the results in FIG. 2. Combined with no sign change for $\varphi\to\varphi+\pi$, it shows a parity violation P=1P=-1 for $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$.
While the components of the contraction of the EM field strength tensor $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$
@@
Appendix C Symmetries of the electric and magnetic fields
@@
$E^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=-E^{r}_{\rm cos/sin}\left(r,\theta\right)~,\quad B^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=B^{r}_{\rm cos/sin}\left(r,\theta\right)~,$ (72)
@@
which can be clearly seen in FIG. 4-5. We further have
@@
These transformations are equivalent to $\overset{\rightharpoonup}{E}\rightarrow-\overset{\rightharpoonup}{E}$ and $\overset{\rightharpoonup}{B}\rightarrow\overset{\rightharpoonup}{B}$ under $\overset{\rightharpoonup}{r}\rightarrow-\overset{\rightharpoonup}{r}$. Therefore the induced electric field still keep a parity symmetry with P=1P=-1 as a vector, and the induced magnetic field keeps the parity symmetry with $P=+1$ as a pseudo-vector, same as the EM background.